Qd=200-5p Qs=40+30p tax= 5% what will be equilibrium price and quantity
1
Expert's answer
2019-03-01T10:43:08-0500

Qd=200-5p Qs=40+30p tax= 5%.

In equilibrium before tax:

Qd = Qs,

200 - 5p = 40 + 30p,

p = 4.57,

Q = 177 units.

The equilibrium price and quantity after task are:

200 - 5p = 40 + 30*(1 - 0.05)*p,

33.5p = 160,

p = 4.78,

Q = 200 - 5*4.78 = 176 units.